# AsynFusion: Towards Asynchronous Latent Consistency Models for Decoupled Whole-Body Audio-Driven Avatars

### v1

@@ 4.1 Datasets
4.2 Implementation Details
Our experiments are conducted on two NVIDIA A100 (40G) GPUs. For the BEAT dataset, the model was trained for 1,000 epochs using a batch size of 1,600. For the SHOW dataset, to accommodate longer sequences and higher frame rates, training was extended to 1,600 epochs with a batch size of 700. Additionally, for the BEAT2 dataset, a batch size of 128 was used over 800 epochs. We compare AsynFusion with several state-of-the-art methods, placing particular emphasis on DiffSHEG [9] and Combo [60] due to their recent advances in joint expression-gesture generation. For the BEAT dataset, our benchmarks include DiffSHEG [9], which employs a unified diffusion framework with unidirectional information flow, and CaMN [32], an LSTM-based model for multimodal feature fusion. Additionally, we assess gesture synthesis against diffusion-based approaches such as DiffGesture [68], DiffuseStyleGesture [62], and LDA [1]. For the SHOW dataset, our comparisons primarily involve Protalk [37], alongside TalkSHOW [65], EMAGE [33] and Combo [60], which propose alternative architectures for coordinated expression-gesture generation. Finally, for the BEAT2 dataset, we mainly compare our method with DiffSHEG [9], EMAGE [33] and Combo [60]. All models use axis-angle rotation representation and are conditioned on both audio and speaker identity. For gesture-only models (DiffGesture, DiffuseStyleGesture, and LDA), we extend their capabilities by training separate models for facial motion. While our focus is on upper body motion, the framework supports full-body animation. Special attention is given to comparing our bidirectional feature interaction and asynchronous sampling with DiffSHEG’s unidirectional approach.
@@ 4.3 Qualitative Evaluation
We present an visualization of our method in Fig. 3, which (a) shows a sequence of generated motions for the speech “Well, Hello There, My name is…”. The visualization demonstrates the effectiveness of our bidirectional feature interaction mechanism in coordinating facial expressions and body gestures. Specifically, during the greeting phrase “Hello There”, our model generates synchronized hand-raising gestures that naturally align with the corresponding facial expressions. The character’s face displays appropriate emotional engagement (transitioning from neutral to friendly expressions) while the hands move in a coherent greeting motion. Such coordination is particularly evident in the natural timing between mouth movements and hand gestures, where the peak of the greeting gesture coincides with the emphasized part of the speech. (b) highlighting motion generation for the speech segment “And I always tell…”. The sequence demonstrates the superior coordination capabilities of AsynFusion compared to baseline methods. In our results, as the speaker emphasizes the word “always”, the model generates a progressive hand-raising gesture that aligns naturally with the speech intensity, indicated by the ascending red line in the figure. This dynamic gesture is synchronized with appropriate facial expressions that transition from neutral to emphatic, creating a cohesive non-verbal emphasis. This example illustrates how the bidirectional information flow enables more natural and coordinated motion synthesis compared to previous approaches that handle facial expressions and body gestures separately or with only unidirectional information flow.
@@
BEAT2 [33] CaMN [32] - - - 6.644 6.769 10.86
DSG [62] - - - 8.811 7.241 11.49
TalkSHOW [65] 7.594 - - 6.209 6.947 13.47
EMAGE [33] 7.310 - - 5.512 7.724 13.06
DiffSHEG[9] 8.329 - - 10.51 5.55 10.91
ours 7.128 - - 5.436 7.612 14.23
Table 1: Quantitative comparison and ablation study on BEAT [30], SHOW [65] and BEAT2 [33] datasets. Best results in each category are in bold; second best are underlined.
@@ 4.4 Quantitative Evaluation
We evaluate our method using a comprehensive set of metrics, including Fréchet distances (FMD, FGD, FED) to assess overall, gesture, and expression quality, along with diversity (Div), beat alignment (BA). These metrics offer a robust evaluation of distributional and temporal aspects of the generated motions. As shown in Table 1, AsynFusion consistently surpasses existing methods on both the BEAT, SHOW and BEAT2 datasets. On BEAT, our model achieves the lowest FMD (312.46) and FED (316.97), outperforming DiffSHEG (324.67 FMD, 331.72 FED) and other methods. AsynFusion also records the best FGD (421.58), producing smoother and more stable gestures than CaMN (1635.44) and diffusion-based approaches like DiffGesture (23700.91) and DSG (1907.58), which exhibit more jitter. In terms of motion dynamics, AsynFusion strikes a better balance between smoothness and expressiveness. Our model generates more dynamic gestures (BA = 0.917) compared to CaMN (BA = 0.793), while closely matching ground truth levels (BA = 0.915). Gesture diversity (Div = 0.561) also improves upon DiffSHEG (Div = 0.536), approaching real data diversity (Div = 0.819). On SHOW, AsynFusion achieves the lowest FMD (3.098), and the highest gesture diversity (Div = 12.53), outperforming Combo(Div = 10.36), TalkSHOW (Div = 12.40) and Probotalk (Div = 10.45). Our model also leads in beat alignment (BA = 0.8701) and gesture stability (FGD = 2.049), outperforming TalkSHOW and Combo. On BETA2, AsynFusion achieves the lowest FMD (7.128) and FGD (5.436), outperforming EMAGE (FMD = 7.310 and FGD = 5.512), and the highest gesture diversity (Div = 14.23).
@@ 5 Ablation Study
In this section, we conduct comprehensive ablation studies to validate the effectiveness of AsynFusion’s key components and design choices on BEAT Dataset [32]. Specifically, we examine (1) the impact of different feature interaction strategies in our Dual-branch DiT Architecture, (2) the efficiency of our asynchronous sampling approach compared to synchronized alternatives, and (3) the computational advantages of our method versus existing approaches.
Impact of Feature Interaction Strategies. To assess the optimal interaction mechanism between the expression and gesture branches, we compare five variants: (1) No Interaction, where both branches operate independently without information exchange; (2) Unidirectional Flow ($E\rightarrow G$), which allows only expressions to influence gestures (similar to DiffSHEG); (3) Unidirectional Flow ($G\rightarrow E$), where only gestures affect expressions; (4) Naïve Fusion, which simply concatenates expression and gesture features in a single diffusion transformer block; and (5) our Bidirectional Interaction design. As shown in Table 2, the Bidirectional Interaction strategy significantly outperforms all alternatives across evaluation metrics. The No Interaction baseline demonstrates poor coordination between expressions and gestures (FMD = 352.41), confirming the necessity of cross-branch communication. While both unidirectional variants show improvements over the baseline, they exhibit imbalanced performance—expression-to-gesture flow excels in facial metrics but underperforms in gesture dynamics (FGD = 435.24), while gesture-to-expression flow shows the opposite trend. This finding is consistent with the conclusions in DiffSHEG. The model with a unidirectional flow from gesture to expression even performed worse than the results of naive fusion approach. This observation inspired our design of asynchronous LCM sampling, as we discovered that the facial component requires significantly fewer LCM sampling steps to achieve competitive results. The Naïve Fusion approach, despite allowing information exchange, fails to effectively model the nuanced relationships between modalities (FMD = 340.16). In contrast, our Bidirectional Interaction design achieves the best performance (FMD = 312.46, FED = 316.97, FGD = 421.58), demonstrating the importance of dynamic, balanced information flow between expressions and gestures.
Efficiency of Asynchronous LCM Sampling. We next evaluate our sampling strategy by comparing three approaches: (1) Traditional DDIM sampling with 25 steps for both branches, (2) Synchronized LCM sampling with 8 steps for both branches, and (3) our Asynchronous LCM approach with 4 steps for the expression branch and 8 steps for the gesture branch. As shown in Table 3, DDIM sampling achieves high quality but at substantial computational cost (56.4s per sequence). The Synchronized LCM approach reduces inference time by 67% (18.6s) with only a marginal quality decrease (FMD from 312.46 to 318.13). Our Asynchronous LCM strategy further improves efficiency (15.9s) by recognizing that facial expressions converge more rapidly than body gestures. Notably, this adaptive approach maintains competitive quality (FMD = 320.59) while reducing inference time by 72% compared to DDIM. The results demonstrate that our asynchronous strategy effectively balances quality and computational efficiency by adapting the sampling process to the inherent characteristics of each motion component.
@@
Table 3: Comparison of different sampling strategies.
Method Inference Time (s) GFLOPs
EMAGE [33] 20.3 18.3
DiffSHEG [9] 28.6 24.7
ProbTalk [37] 5.1 9.7
Our(Sync LCM) 18.6 12.5
Our(Async LCM) 15.9 10.6
Table 4: Comparison of computational efficiency.
Computational Efficiency Comparison. Finally, we compare AsynFusion’s computational efficiency against state-of-the-art methods for co-speech motion synthesis. As illustrated in Table 4, our method achieves the fastest inference time (15.9s per sequence) and lowest computational complexity (58.4 GFLOPs) while maintaining competitive or superior quality metrics. Compared to unified approaches like Probtalk, AsynFusion reduces inference time by 62% and computational requirements by 57%. Even against recent efficient methods like DiffSHEG, our approach shows a 43% reduction in inference time and 38% decrease in GFLOPs. This significant improvement stems from our dual-branch architecture and asynchronous sampling strategy, which effectively parallelize computation and optimize sampling steps for each motion component. These results confirm that AsynFusion successfully addresses the computational efficiency challenges in audio-driven avatar animation while preserving high-quality motion synthesis.
@@ 6 Conclusion
In this paper, we present AsynFusion, a novel framework for audio-driven avatar animation that advances joint expression-gesture generation. Our key contributions include a dual-branch DiT architecture with bidirectional feature interaction, enabled by the CoSync module, which improves the coordination between facial expressions and body gestures. Additionally, we introduce an asynchronous sampling strategy using Latent Consistency Models, significantly reducing computational overhead while maintaining high-quality motion synthesis.

### v2

@@ 4.4 Qualitative Evaluation
Fig.  3 presents visualizations of motions generated by our method. In (a), for the speech “Well, Hello There, My name is…”, our model produces synchronized hand-raising gestures and facial expressions, demonstrating the effectiveness of bidirectional interaction. Notably, during “Hello There”, the greeting gesture aligns with a smooth transition from neutral to friendly expressions, and gesture peaks match speech emphasis. In (b), for “And I always tell…”, AsynFusion outperforms baselines by generating a rising hand gesture synchronized with the emphasis on “always”, as indicated by the red intensity curve. The facial expression shifts accordingly, showcasing cohesive non-verbal emphasis enabled by our bidirectional design.
Table 1: Quantitative comparison and ablation study on BEAT [19], SHOW [37] datasets. Best results in each category are in bold; second best are underlined.
@@ 4.5 Quantitative Evaluation
We evaluate our method using FMD, FGD, FED, diversity (Div), and beat alignment (BA) to capture motion quality, expressiveness, and temporal alignment. As shown in Table 1, AsynFusion consistently outperforms prior methods on both BEAT and SHOW datasets. On BEAT, it achieves the best FMD (312.46), FED (316.97), and FGD (421.58), indicating superior overall quality and gesture stability. It also improves gesture diversity (Div = 0.561) and beat alignment (BA = 0.917), closely matching ground truth. On SHOW, AsynFusion achieves top scores in FMD (3.098), Div (12.53), BA (0.8701), and FGD (2.049), surpassing Combo, TalkSHOW, and Probotalk across all metrics.
@@ 5 Ablation Study
In this section, we conduct comprehensive ablation studies to validate the effectiveness of AsynFusion’s key components and design choices on BEAT Dataset [21]. Specifically, we examine (1) the impact of different feature interaction strategies in our Dual-branch DiT Architecture, (2) the efficiency of our asynchronous sampling approach compared to synchronized alternatives.
[…]
Impact of Different Feature Fusion Methods. In our exploration of improving the coordination between facial expressions and body gestures, we explored three distinct feature fusion approaches: Cross Attention (CA), Feature Concatenation (FC), Gated Fusion (GF). Our AsynFusion framework implements Cross Attention, leveraging its bidirectional mechanism to achieve fine-grained control over expression-gesture information flow, thereby producing more natural and expressive animations. We also examined two alternative methods: Feature Concatenation, which simply concatenates expression and gesture features before feeding them into the next DiT block, and Gated Fusion, which employs learnable weights through a gating mechanism to control the fusion process:
$F_{fused}=\sigma(W_{g}[F_{exp};F_{ges}])\odot F_{exp}+$ (19)
$(1-\sigma(W_{g}[F_{exp};F_{ges}]))\odot F_{ges}$
where $\sigma$ represents the sigmoid function, and $W_{g}$ are the learnable weights. As shown in Table 3 This approach adaptively controls each modality’s influence through learnable weights. While Gated Fusion achieves better modality balance than simple concatenation, it lacks the sophisticated bidirectional interaction of Cross Attention, making it less capable of capturing the subtle dependencies in natural human behavior.
Table 3: Ablation study on different Fusion Strategy. Best results in each category are in bold; second best are underlined.
Fusion Strategy Holistic Expression Gesture
FMD $\downarrow$ FED $\downarrow$ FGD $\downarrow$
Feature Concat 340.16 340.23 467.33
Gated Fusion 325.46 329.15 443.74
Cross Attention 312.46 316.97 421.58
Efficiency of Asynchronous LCM Sampling. We evaluate three sampling strategies: (1) DDIM (25 steps), (2) Synchronized LCM (8 steps), and (3) our Asynchronous LCM (4 steps for expression, 8 for gesture). As shown in Table 4, DDIM achieves the best quality (FMD = 312.46) but is slow (56.4s). Synchronized LCM reduces time by 67% (18.6s) with minimal quality drop (FMD = 318.13). Our Asynchronous LCM further improves efficiency (15.9s, 72% faster than DDIM) while maintaining competitive quality (FMD = 320.59), showing the benefit of adapting sampling to the convergence speed of each branch.
Table 4: Comparison of different sampling strategies.
@@ 6 Conclusion
A key limitation of AsynFusion lies in its dependency on training data quality. Our model, like other deep learning approaches, inherits both desired behaviors and undesirable artifacts from the training datasets. For example, when trained on BEAT [21], the generated motions exhibit jittering artifacts similar to those in the dataset’s female character animations. Similarly, expression discontinuities from the SHOW dataset appear in our synthesized results. These observations highlight that future improvements in motion synthesis may rely as much on better data collection and cleaning as on model architecture advances. Future research could focus on end-to-end multimodal foundation models for motion synthesis, processing speech, text, and video simultaneously. Large-scale pretraining across diverse data sources could enable deeper understanding of verbal and non-verbal communication patterns. This approach could enhance motion diversity and naturality through universal representations, while advanced cross-modal pretraining could improve the capture of speech-motion correlations for more nuanced animations.
Acknowledgement
This work was supported by the National Natural Science Foundation of China under Grant No. 62441617 and 62476224. It was supported by the Postdoctoral Fellowship Program and China Postdoctoral Science Foundation under Grant No. 2024M764093 and Grant No. BX20250485, the Beijing Natural Science Foundation under Grant No. 4254100, the Fundamental Research Funds for the Central Universities under Grant No. KG16336301, and by Beijing Advanced Innovation Center for Future Blockchain and Privacy Computing.